OpenStudy (anonymous):

What is the sum of the arithmetic sequence 136, 124, 112 …, if there are 36 terms? –2,736 –2,664 –2,628 –2,700

OpenStudy (anonymous):

I believe it is -2664?

OpenStudy (mr.math):

Are you sure?

OpenStudy (mr.math):

What's the common difference between each successive members?

OpenStudy (mr.math):

Oh sorry, I read the choices as the members of the sequence. My bad!

OpenStudy (ash2326):

a1=136 an=a1+(n-1)d an=136+(36-1)(-12) an=136-420 an=-284 an=-284 is the 36th term sum =n/2(a1+a36) =36/2(136-284) =18*(-148) -2664

OpenStudy (mr.math):

-2664 is right.

OpenStudy (anonymous):

Haha thank you :)

OpenStudy (mr.math):

Haha you're welcome :P